OpenStudy (anonymous):

(m2-8m+16)/(5m2+m) can you please factor the denominator and excludes values for m. Please explain step by step.

OpenStudy (tkhunny):

You should first remember your Order of Operations. You have written m2-8m+16/5m2+m = \(x^{2} - 8m + \dfrac{16}{5m^{2}} + m\) I'm guessing you intended to write (m2-8m+16)/(5m2+m) = \(\dfrac{m^{2}-8m+16}{5m^{2}+m}\) Have you considered factoring both numerator and denominator?

OpenStudy (anonymous):

Refer to the attached Mathematica presentation.

OpenStudy (anonymous):

Ok. Thank you. Can you show me how you arrived at -1/5. I do not understand your graph. If you factor the denominator (5m2+m) you get 5m(m+1). The professor shows a example (5-9w)/(9w2-4) 9w2-4 =0 Difference in square which I can factor (3w-2)(3w+2)=0 Set each factor equal to zero 3w-2=0 or 3w+2 = 0 Add or subtract 2 from both sides 3w=2 or 3w=-2 Divide both sides by 3 w=2/3 or w =-2/3 These are the excluded values to my expression